OpenStudy (anonymous):

Find the slope: (9,2) and (5,8)

OpenStudy (anonymous):

m= -3/2

OpenStudy (anonymous):

\[8-2/5-9= -3/2\]

OpenStudy (anonymous):

\[(2-8)/(9-5)=(-6)/4=-3/2\]

OpenStudy (anonymous):

thanks